# Automatic Image Colorization with Convolutional Neural Networks and Generative Adversarial Networks

### v1

@@
equal contribution, ranked in alphabetic order
@@ 2 Related Works
[5] takes the lead in investigating fully-automatic colorization with deep neural networks (DNNs) and formulates image colorization as a regression problem: the DNN takes the extracted feature descriptors at each pixel as input and outputs the continuous values of the colored channels at the corresponding pixel, with the loss function to be the Mean Square Error (MSE) between the predicted colored channel values and the ground truth. [20] first formulates the task as a classification problem by quantizing the $ab$ channels of the CIE $Lab$ color space into 313 discrete $ab$ pairs and using the multinomial cross entropy loss to train the model. [10, 15] uses conditional Generative Adversarial Networks (cGANs) for automatic colorization of grayscale images.
@@ 3.2 GAN-based Colorization
To better control the color space, we separate the brightness channel and color channels using the CIE $Lab$ color space, where we only need to predict two color channels $ab$ in the generator.
@@ 4.1 Evaluation Metrics
Besides, we use peak-signal-to-noise ratio (PSNR) and structrural similarity index (SSIM index) to measure image qualities [9]. PSNR is defined in log form of the mean sqaure error and we calculate PSNR for each generated images( after transformation from $Lab$ to rgbrgb). SSIM is a method to measure the similarity between images. We calculate average PSNR and SSIM values on the test dataset and the results are shown below.

### v2

@@
All authors have equal authorship and equal contribution, ranked in alphabetic order. First version of this paper was completed in 2021.
@@ 2 Related Works
[5] takes the lead in investigating fully-automatic colorization with deep neural networks (DNNs) and formulates image colorization as a regression problem: the DNN takes the extracted feature descriptors at each pixel as input and outputs the continuous values of the colored channels at the corresponding pixel, with the loss function to be the Mean Square Error (MSE) between the predicted colored channel values and the ground truth. [20] first formulates the task as a classification problem by quantizing the abab channels of the CIE LabLab color space into 313 discrete abab pairs and using the multinomial cross entropy loss to train the model. [10, 15] uses conditional Generative Adversarial Networks (cGANs) for automatic colorization of grayscale images.
@@ 3.2 GAN-based Colorization
To better control the color space, we separate the brightness channel and color channels using the CIE LabLab color space, where we only need to predict two color channels abab in the generator.
@@ 4.1 Evaluation Metrics
Besides, we use peak-signal-to-noise ratio (PSNR) and structrural similarity index (SSIM index) to measure image qualities [9]. PSNR is defined in log form of the mean sqaure error and we calculate PSNR for each generated images( after transformation from LabLab to rgbrgb). SSIM is a method to measure the similarity between images. We calculate average PSNR and SSIM values on the test dataset and the results are shown below.